# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 2. Legal, Ethical and Statistical Confidentiality Requirements
We present an example here that foreshadows the results of our study. Suppose a user of published census data wishes to learn the racial and ethnic makeup of each individual in a small neighborhood, for example, a census block. The census block contains 12 non-Hispanic Whites, 5 non-Hispanic Blacks, 1 non-Hispanic Asian, 1 Hispanic White and 1 Hispanic Black. The analyst also has other census block-level information on age and sex sufficient to perform the reconstruction studied in this paper. For each individual, the analyst might guess “non-Hispanic White” (the modal value) or guess in proportion to the observed frequencies. Other forecast models are feasible (e.g., guess each of the 126 possible race and ethnicity values used in the schema for the 2010 Census with probability $1/126$ or guess using a model that combines data at different levels of geography), but this example can be adapted if such models are used. The modal or proportional block-level forecast may be a statistical or generalizable inference, which is permitted by the legislation and ethical standards governing statistical agencies. [^2] Formally, whether such an inference is statistical or nonstatistical depends on how much other data are also released about the same individuals. As the rest of the example makes clear, the ensemble of published data can enable nonstatistical inferences even in cases where the use of block-level race and ethnicity data by themselves might not. Now suppose the user independently knows the name, sex and age of each person in this block. The same user takes published tables from the census and creates 20 records with values of sex, age, race, and ethnicity that are consistent with the information in the block-level tables (and possibly tract- and county-level tables containing this block). The user then associates the race and ethnicity from these reconstructed census records with the 20 persons in the block by matching on sex and age. Now that user has name, sex, age, race, and ethnicity for every person on the block, coded consistently with the schemas used in the published tables.
One measure of the data user’s gain from the reconstructed microdata relative to having only the race by ethnicity counts for the census block is the increased precision of the race and ethnicity forecast for each person compared to the precision possible if no data for that person were used in the published census results. Deleting the record of a non-Hispanic White, in this case, has an ambiguous effect on the precision of such race/ethnicity forecasts. If the user adopted the modal prediction method, the mode is unchanged by deleting one non-Hispanic White, and so the forecast precision is the same with or without that record. If the user adopted the proportional guess method, then deleting one non-Hispanic White record, slightly lowers the proportion of non-Hispanic Whites; hence, there is a slight reduction in precision when the record is removed from the census data. Now consider the sole Hispanic White on the block. If that person’s record is removed, there is no possibility that the forecast models for race/ethnicity considered here can get the right combination. Both modal guessing and proportional guessing have no chance of being correct; that is, they have precision of exactly zero in the absence of the Hispanic White’s census data. On the other hand, if the user’s independent data on age and sex for the Hispanic White are close enough to the values in the reconstructed census data, then the user will correctly infer Hispanic White for this person when data from all 20 persons on the block are used in the census tables. There is an infinite gain in prediction accuracy when the record for the unique Hispanic White is included in the published data. The user’s externally supplied name, sex, and age of the target were associated with a report of Hispanic White on the census record. When reliably correct, that is a prohibited, nonstatistical inference because it is only possible when the target person’s record is used to produce the published tables. [^3] In traditional statistical disclosure limitation, e.g., [duncan:etal:2011:statistical, p. 30], the inference associating the Hispanic White race/ethnicity with a specific, named person in the census block might be called either an identity or attribute disclosure—both of which are prohibited—depending on the setup of the attack. See [garfinkel:near:etal:2023, ] and [kifer:etal:2022, ].
@@
One might ask “What’s the harm?” That’s a perfectly legitimate question, even if outside the scope of the legal requirements governing the Census Bureau. One can see the harm by considering two routine uses of census data: redistricting and local demographics. Experts in both fields maintain databases that contain names, addresses, and some demographic data. They routinely update these databases. Redistricting experts use voter registration lists and purchase commercial data. Demographers use school district and commercial data. Both groups have mission-valid reasons to improve the accuracy of those databases. Both groups do this using their models and microdata-level conversions of published census tables [jarmin:etal:2023]. When those census tables permit nonstatistical (non-LOO) inferences, these users gain access to information about a specific person that is only possible because that person responded to the census and that response was used in the tabulations. Even though redistricters have legitimate interest in these data, a data steward who has made a confidentiality pledge to collect the data should not subsequently violate that pledge by permitting uses that depend specifically on the response provided. Individuals have a right to privacy that is reinforced by the same statute that ensures the confidentiality of the response should the individual answer the census. That right extends to protecting their responses from use by redistrictors or school districts via privacy-violating inferences. Such protection is the point of data confidentiality laws—they balance the utility of published data against the potential for privacy breaches. A user might want to know a characteristic more accurately than a statistical (LOO) inference permits, but the data steward should not facilitate that learning by publishing data that permit strong non-LOO inferences. To prevent the extraction and nonstatistical use of personal information, statistical agencies must periodically analyze their SDL methods because nonstatistical actors, and especially malicious actors, do not publicly advertise their plans or methods. The goal of this paper is to assess the effectiveness of the SDL methods used in the 2010 Census in preventing nonstatistical inferences based on the published data.
We analyze the effectiveness of the confidentiality protections applied to the data products published from the 2010 Census by simulating what we call a reconstruction-abetted reidentification attack. We first attempt to reconstruct the underlying microdata for the features of census block, sex, age, race, and ethnicity from publicly released tables. We then attempt to match those reconstructed records to (a) records in low-quality commercial databases acquired during the conduct of the 2010 Census with person and address identifiers (representing attackers with the same quality contemporaneously available information in 2010) and (b) records containing a limited subset of variables from the 2010 Census itself (person identifier, census block, sex and age; representing attackers higher-quality contemporaneous external information). We used shared variables, called “key variables” in [duncan:etal:2011:statistical, p. 20] or “quasi-identifiers” in [garfinkel:near:etal:2023, p. 49], to determine how accurately we could link records and then infer the race and ethnicity of the persons in (a) the commercial records and (b) the quasi-identifier-only version of the 2010 Census records. We compare the results in several scenarios against baseline attackers who predicted using either (a) the most common race and ethnicity pair in the block (modal prediction from public tables) or (b) the race and ethnicity pair proportional to the distribution of block-level race by ethnicity combinations (proportional prediction from public tables). In all cases, we assess the prediction accuracy using the full set of 2010 Census variables (person identifier, census block, sex, age, race and ethnicity). For census block, sex, race, and ethnicity we always used the same schema. For age, we considered two different schemas: the block-level table schema (38 narrow age categories) and the tract-level schema (111 exact age categories). For the person identifier, we used the same vintage of the Census Bureau’s production record linkage system to replace the reported name and address with internal identifiers routinely used for person and household record linkage.
@@ 2.1. Why Protections Are Required
Title 13 of the U.S. Code mandates that information gathered from individuals and establishments remain confidential. Specifically, 13 U.S. Code § 8(b) allows the Census Bureau to “furnish copies of tabulations and other statistical materials which do not disclose the information reported by, or on behalf of, any particular respondent,” and § 9 prohibits the release of “any publication whereby the data furnished by any particular establishment or individual under this title can be identified.” First and foremost, the Census Bureau is required to protect the confidentiality of respondents by law. Additionally, it is in the best interest of data quality that the public trust the Census Bureau to protect their data so that truthful responses are given, especially to sensitive questions [Childs:Development:JSM:2012, childs:etal:2019, Childs:Confidence:SP:2015, childs:et:al:2020:bigsurv].
@@
  • Age, sex, race, and ethnicity data about children are often missing in commercial databases due to legal restrictions, because information about children is generally considered more sensitive.
  • Household composition may be a sensitive subject in some areas and the decision to reveal this information in identifiable form should be up to the household and not the Census Bureau according to the principles guiding statistical agencies [nasem:2021, p. 3]. This includes the detailed (census block-level) location of same-sex spouses, unmarried partners, mixed-race households, households with adopted children, older individuals living alone, etc. Thus, to encourage accurate reporting, the Census Bureau should protect the confidentiality of those responses.
@@
Another argument that some give against strong confidentiality protections for census data is that there is so much personal information data “out there” that the census data does not pose an incremental risk. While there are large amounts of data available externally, the accuracy of this information is generally unknown. Our experiments demonstrate that circa 2010 external data were indeed inaccurate, or at least very noisy, compared the the decennial census data; however, circa 2020 external data are much more accurate [brown:etal:2023]. Additionally, if a data steward adopts the policy that data they collect should not be protected because it is already “out there,” then survey response rates would drop: “why should I fill out the survey if my data is already out there, just use that and don’t bother me?” But even when this position is accepted by a statistical agency, the relevant confidentiality statutes still require that the census responses, including any generated from external data, be protected. One might be able to learn respondent-specific information collected on the census from other sources, but the census publications cannot facilitate this learning [statcan-confidentiality-admin-data].
@@ 2.2. Household Data Swapping in the 2010 Census
In the 2010 Census, the agency used noise infusion via targeted data swapping as the primary SDL framework. Households deemed at high risk for reidentification were swapped with a higher probability, but all records that were not entirely imputed had some chance of being swapped [mckenna2018disclosure]. High-risk households were those in low-population census blocks or those who had a member with a unique race category in the census block. Pairs of swapped households matched on two key demographic variables: the total number of persons and the number of adults living in the household. Once swap pairs were determined, the geographic identifiers were swapped, effectively relocating the two records in different geographies from their 2010 Census values. The swapped file was used to produce all tabular and microdata products. Data swapping, by itself, is highly susceptible to attack [kifer:10.1145/2783258.2783369]. For example, if the swap rate is 1%, then each record has a 99% chance of being unaltered and hence an attacker linking to a record can have high confidence that attributes learned from the record are correct. Furthermore, external data can help identify swapped records and even undo data swaps. If a household A, placed in a different census block in the swapped data than its census response, can be linked to external data based on matching attributes that are not affected by the swap, but does not match based on some of the attributes changed in the swap, then that household was likely targeted for swapping. Furthermore, if a household B in a nearby census block is a better match on the attributes affected by swapping, then it is likely that A and B were swapped with each other. For this reason, swapping is often paired with aggregation in an attempt to further thwart attackers. Note that our reconstruction-abetted reidentification attack only attempts to undo the aggregation and does not take the further step of undoing the swapping protections.
@@ 2.3. Related Work
Reconstruction and reidentification attacks have been studied both theoretically and practically. [Dinur:Nissim:2003:RIW:773153.773173, ] demonstrated conditions under which a database could be reconstructed even when only perturbed queries were reported. [dworkSteinkeUllman2017, ] provided an overview of reidentification and reconstruction attacks. Notable reconstruction and reidentification attacks include reidentification of individuals in a homicide database by making comparisons with public social security data [ochoa2001reidentification], reidentification of patients in de-identified pharmaceutical marketing data using publicly available hospital discharge and ambulatory claims data and voting list data [sweeney2011patient], an attack against official foreign-trade statistics released in Brazil that reidentified companies performing import transactions [favato2022novel], a genomics data reconstruction attack [ayoz2021genome], and reidentification of individuals in census microdata publications with very high precision [rocher:et:al:2019]. [dick:etal:2023, ] demonstrated a simulation-based reconstruction attack on census data with a very high success rate in identifying population uniques. There are many more examples in [Garfinkel:Deidentification:NIST:2015, ] and [garfinkel:near:etal:2023, ].
In our results we include two separate statistical baseline models for comparisons that reflect a random-guessing attacker using statistical (LOO) inference to assign race and ethnicity. We find that the reconstruction aids the attacker substantially in correctly inferring the race and ethnicity of reidentified persons whose race and ethnicity differ from the majority race and ethnicity in the census block where they live. Specifically, we find that for a large set of reconstructed records, the attacker can prove that they have found the exact image of the source census record in the reconstruction using only public data, that this image is unique in the population, and that the race and ethnicity pair associated with the record is the same as the pair on the confidential census record even when that person is the only member of the race and ethnicity group in their census block. In other words, the only reason that the attacker found the correct race and ethnicity pair was because that person’s response was used to produce the underlying tables that supported the reconstruction. This is a prohibited nonstatistical use of census data made possible by the use in the 2010 Census of SDL methods that did not anticipate reconstruction-abetted reidentification attacks.
2.4. Statistical Inference vs. Breaches of Confidentiality
@@
Privacy and confidentiality protections are subtle concepts that give rise to many misunderstandings. It is a common but incorrect belief that breaches of confidentiality, colloquially known as privacy breaches, occur when a dataset is used to make any harmful or unwanted inference about an individual. Such an error has made its way into many peer-reviewed papers [jarmin:etal:2023, SI section 5]. In order to see where the problem lies, we first discuss the canonical “smoking causes cancer” thought experiment [dwork:2011] and then discuss possible confidentiality concerns in the 2020 Census data. A more complete version of these arguments can be found in [kifer:etal:2022, ].
@@
The first Cancer Prevention Study, also known as CPS-I, followed a cohort of volunteers from 1959 to 1972 and conclusively established the link between smoking tobacco cigarettes as a cause of death from lung cancer and coronary heart disease [cps1]. As a result of this study, we know that persons who smoke have a much higher risk of developing lung cancer. Such inferences about smokers can definitely be unwanted, as they result in higher health and life insurance premiums. Persons born after 1972 may be subject to this inference caused by the study; however, since their data were not used for the study (they were not born until after it was completed), the study cannot possibly be considered a privacy breach of their data. For those people, one would say that the inference is purely statistical in nature. Another way to phrase this is that the link between smoking and lung cancer is a population property, or statistical use (in the sense of the 2018 Confidential Information Protection and Statistical Efficiency Act, Title III of the 2018 Foundations of Evidence-based Policymaking Act; 44 U.S. Code § 3561(12)) that was uncovered with the help of the data set. This is exactly why data sets are collected and published.
@@
In contrast, an unwanted inference is a privacy breach when it is specifically caused by the inclusion of the individual’s information in the dataset from which the inference was made, what we call non-LOO inference. Now consider a hypothetical CPS-I study participant named Charlie who was a lifelong smoker. Suppose that as a result of the study, Charlie’s insurance company decided to ask enrollees whether or not they smoked and charge a higher premium if they did. As a result, Charlie was harmed by the result of the study in which he was a participant. Is this now a privacy breach? To answer that question, we turn to causal reasoning, and specifically consider a counterfactual in which Charlie had not participated in the study; that is, we compare the non-LOO inference with a properly computed LOO one. Would the outcome of the study have been different enough without Charlie’s participation to change the findings—thus changing whether the insurance company enacted their policy of charging smokers a higher premium policy which harmed Charlie? Given the strength of the findings in CPS-I [hammond:horn:1954], we can be confident that the study would have drawn the same conclusions even if Charlie had not participated. Therefore, this example would also not be considered a confidentiality breach.
@@
One also must be careful about the conflation of harmful inferences with privacy violations for another reason, which is also related to privacy and trust in statistical agencies, again as they are expressed in 44 U.S. Code § 3561-4 and [nasem:2021, ]. Specifically, sometimes a statistical inference should not be allowed even if it is not privacy violating. A statistical agency might be asked to produce data on a particular sensitive population that could reasonably expect harm from those data even if they passed the LOO inference test. For example, in 2010, communities with a high percentage of same-sex married couples in states that did not permit such weddings might expect harm even if only a statistical summary were published. A statistical summary that could have affected the 2020 Census is the proportion of residents in the census block who were not U.S. citizens. Whether or not the government has a legitimate statistical interest in these populations is indeed a policy question, but the policy concerns ingesting the data with the intention to publish summaries in the first place. Barring the collection or publication of data on the grounds that even statistical inference may be harmful is a policy concern. In this paper, we presume that the agency’s legitimate interest in supporting statistical inference has already been determined; that is, the collection of data in the decennial census and their publication for statistical purposes are authorized by Congress and undertaken consistent with the required trust and confidentiality policies.
In practice, the distinction between privacy-violating and statistical inferences is not always so clear-cut as in our smoking example. Information obtained from individuals is often aggregated, fields are suppressed (e.g., direct identifiers), and some SDL framework beyond aggregation may also be used. Still, it is important to disentangle what could be learned from statistical inference versus what can be specifically learned or caused by a person’s participation/inclusion in a study or dataset. This idea is the basis for research in the scientific field of “differential privacy” [Dwork:Algorithmic:Book:2014].
3. 2010 Census and Commercial Data Inputs

### v2

@@ 2.1. Defining and diagnosing confidentiality violation in 2010 SDL methods
We present an example here that foreshadows the results of our study. Suppose a user of published census data wishes to learn the racial and ethnic makeup of each individual in a small neighborhood, for example, a census block. The census block contains 20 persons: 12 non-Hispanic Whites, 5 non-Hispanic Blacks, 1 non-Hispanic Asian, 1 Hispanic White and 1 Hispanic Black. Suppose also that the user has other census block-level information on age and sex sufficient to perform the reconstruction studied in this paper, and that all individuals are 20-year-old males, except for the non-Hispanic White person, who is a 21-year-old male. For each individual, the user might guess “non-Hispanic White” (the modal value) or guess in proportion to the observed frequencies. Other forecast models are feasible (e.g., guess each of the 126 possible race and ethnicity values used in the schema for the 2010 Census with probability $1/126$ or guess using a model that combines data at different levels of geography), but this example can be adapted if such models are used. The modal or proportional block-level forecast may be a statistical or generalizable inference, which is permitted by the legislation and ethical standards governing statistical agencies. [^3] Formally, whether such an inference is statistical or nonstatistical depends on how much other data are also released about the same individuals. As the rest of the example makes clear, the ensemble of published data can enable nonstatistical inferences even in cases where the use of block-level race and ethnicity data by themselves might not. Now suppose the user independently knows the name, sex and age of each person in this block. The same user takes published tables from the census and creates 20 records with values of sex, age, race, and ethnicity that are consistent with the information in the block-level tables (and possibly tract- and county-level tables containing this block). The user then associates the race and ethnicity from these reconstructed census records with the 20 persons in the block by matching on sex and age. Now that user has name, sex, age, race, and ethnicity for every person on the block, coded consistently with the schemas used in the published tables.
One measure of the data user’s gain from the reconstructed microdata relative to having only the race by ethnicity counts for the census block is the increased precision of the race and ethnicity forecast for each person compared to the precision possible if no data for that person were used in the published census results. Deleting the record of a non-Hispanic White person has a relatively small effect on any of the three kinds of inference: the modal race and ethnicity does not change, so the modal guesser makes the same prediction, the proportional guesser has a slightly smaller chance of guessing non-Hispanic White, and the reconstruction-abetted reidentification attack is slightly less likely to link any given 20-year-old male auxiliary record to a non-Hispanic white reconstructed record. However, for the sole 21-year-old male Hispanic White record, their inclusion or exclusion from the data set is decisive: the modal guesser will still never get this person’s race and ethnicity right, and the proportional guesser will move from a 1/20 probability of guessing Hispanic White to a 0 probability, but the reconstruction-abetted reidentification attack changes from correctly inferring this person’s race and ethnicity with probability 1 to probability 0. Thus, even if we conservatively think of the modal and proportional guessers as always making inferences that are statistical or generalizable [^4] Strictly speaking, an attacker should specify a single method of inference, not multiple; e.g., the attacker could specify a rule for how they will use public information to determine whether to make a modal or proportional guess, or to instead use the reconstruction-abetted reidentification inference. Here, we treat the modal and proportional guessers as if the attacker can try all three methods of attack and keep the one that works best, but this is not possible in practice. This is another sense in which our approach is conservative., the reconstruction-abetted reidentification attack exhibits a much larger change in inferential success than these two baseline attacks. It is this much larger change, caused specifically by the presence or absence of the target person’s record, that we regard as clear evidence of a non-generalizable, confidentiality-violating inference. These inferences can only be highly successful because they have isolated information unique to a target person, so we further regard these as identifying in the sense of Title 13, the statute governing Census Bureau’s confidentiality requirements, which specifically prohibits the Census Bureau from enabling identification of respondents’ data (see Section 2.2).
@@
One might ask: “What’s the harm?” That’s a perfectly legitimate question, even if outside the scope of the legal requirements governing the Census Bureau. One can see the harm by considering two routine uses of census data: redistricting (the re-drawing of geographic boundaries for legislative districts in each U.S. state) and local demographics. Experts in both fields maintain databases that contain names, addresses, and some demographic data. They routinely update these databases. Redistricting experts use voter registration lists and purchase commercial data. Demographers use school district and commercial data. Both groups have mission-valid reasons to improve the accuracy of those databases. Both groups do this using their models and microdata-level conversions of published census tables (jarmin:etal:2023). When those census tables permit nonstatistical (non-LOO) inferences, these users gain access to information about a specific person that is only possible because that person responded to the census and that response was used in the tabulations. Even though redistricters have legitimate interest in these data, a data steward who has made a confidentiality pledge when collecting the data should not subsequently violate that pledge by permitting uses that depend specifically on the response provided. Individuals have a right to confidentiality that is reinforced by the same statute that ensures the confidentiality of the response should the individual answer the census. That right extends to protecting their responses from use by redistricters or school districts via confidentiality-violating inferences. Such protection is the point of data confidentiality laws—they balance the utility of published data against the potential for confidentiality breaches. A user might want to know a characteristic more accurately than a statistical (LOO) inference permits, but the data steward should not facilitate that learning by publishing data that permit strong non-LOO inferences.
@@ 2.2. Why protections are required
Title 13 of the U.S. Code mandates that information gathered from individuals and establishments remain confidential. Specifically, 13 U.S. Code § 8(b) allows the Census Bureau to “furnish copies of tabulations and other statistical materials which do not disclose the information reported by, or on behalf of, any particular respondent,” and § 9 prohibits the release of “any publication whereby the data furnished by any particular establishment or individual under this title can be identified.” First and foremost, the Census Bureau is required to protect the confidentiality of respondents by law. Additionally, it is in the best interest of data quality that the public trust the Census Bureau to protect their data so that truthful responses are given, especially to sensitive questions (Childs:Development:JSM:2012; childs:etal:2019; Childs:Confidence:SP:2015; childs:et:al:2020:bigsurv).
@@
  • Age, sex, race, and ethnicity data about children are often missing in commercial databases (collected and maintained by private companies) due to legal restrictions, because information about children is generally considered more sensitive.
  • Household composition may be a sensitive subject in some areas and the decision to reveal this information in identifiable form should be up to the household and not the Census Bureau according to the principles guiding statistical agencies (nasem:2021, p. 3). This includes the detailed (census block-level) location of same-sex spouses, unmarried partners, mixed-race households, households with adopted children, older individuals living alone, etc. Thus, to encourage accurate reporting, the Census Bureau should protect the confidentiality of those responses.
@@
Another argument that some give against strong confidentiality protections for census data is that there is so much personal information data “out there” that the census data does not pose an incremental risk. While there are large amounts of data available externally, the accuracy of this information is generally unknown. Our experiments demonstrate that circa 2010 external data were indeed inaccurate, or at least very noisy, compared the the decennial census data; however, circa 2020 external data are much more accurate (brown:etal:2023). Additionally, if a data steward adopts the policy that data they collect should not be protected because it is already “out there,” then survey response rates would drop: “why should I fill out the survey if my data is already out there, just use that and don’t bother me?” But even when this position is accepted by a statistical agency, the relevant confidentiality statutes still require that the census responses, including any generated from external data, be protected. One might be able to learn respondent-specific information collected on the census from other sources, but the census publications cannot facilitate this learning (statcan-confidentiality-admin-data).
@@ 2.3. Household data swapping in the 2010 Census
In the 2010 Census, the agency used noise infusion via targeted data swapping as the primary SDL framework. Households deemed at high risk for reidentification were swapped with a higher probability, but all records that were not entirely imputed had some chance of being swapped (mckenna:2018:disclosure). High-risk households were those in low-population census blocks or those who had a member with a unique race category in the census block. Pairs of swapped households matched on two key demographic variables: the total number of persons and the number of adults living in the household. Once swap pairs were determined, the geographic identifiers were swapped, effectively exchanging the geographic location of the two households. The swapped file was used to produce all tabular and microdata products. Data swapping, by itself, is highly susceptible to attack (kifer:10.1145/2783258.2783369). For example, if the swap rate is 1%, then each record has a 99% chance of being unaltered and hence an attacker linking to a record can have high confidence that attributes learned from the record are correct. Furthermore, external data can help identify swapped records and even undo data swaps. If household A, placed in a different census block in the swapped data than its census response, can be linked to external data based on matching attributes that are not affected by the swap, but does not match based on some of the attributes changed in the swap, then that household was likely targeted for swapping. Furthermore, if household B in a nearby census block is a better match on the attributes affected by swapping, then it is likely that A and B were swapped with each other. For this reason, swapping is often paired with aggregation in an attempt to further thwart attackers. Note that our reconstruction-abetted reidentification attack only attempts to undo the aggregation and does not take the further step of undoing the swapping protections.
@@ 2.4. Related work
Reconstruction and reidentification attacks have been studied both theoretically and practically. Dinur:Nissim:2003:RIW:773153.773173 demonstrated conditions under which a database could be reconstructed even when only perturbed queries were reported. dworkSteinkeUllman2017 provided an overview of reidentification and reconstruction attacks. Notable reconstruction and reidentification attacks include reidentification of individuals in a homicide database by making comparisons with public social security data (ochoa2001reidentification), reidentification of patients in de-identified pharmaceutical marketing data using publicly available hospital discharge and ambulatory claims data and voting list data (sweeney2011patient), an attack against official foreign-trade statistics released in Brazil that reidentified companies performing import transactions (favato2022novel), a genomics data reconstruction attack (ayoz2021genome), and reidentification of individuals in census microdata publications with very high precision (rocher:et:al:2019). dick:etal:2023 demonstrated a simulation-based reconstruction attack on census data with a very high success rate in identifying population uniques. There are many more examples in Garfinkel:Deidentification:NIST:2015 and garfinkel:near:etal:2023.
The present work differs significantly from the approaches of (rocher:et:al:2019) and dick:etal:2023. rocher:et:al:2019 were primarily concerned with estimating population uniqueness given sample uniqueness, while we work with a census, where there is no sampling variability. dick:etal:2023 investigated a gradient-descent-based approach to reconstruction, while our approach is based on techniques for mixed-integer linear programming. Neither rocher:et:al:2019 nor dick:etal:2023 directly evaluated the success of reidentification in a decennial census, and neither presented a bound on how different two possible reconstructions could be from one another, though dick:etal:2023 suggested an interesting heuristic measure of confidence in individual reconstructed records. dick:etal:2023 also specifically cite the reconstruction-reidentification attack we discuss here as motivation for their work. Lastly, neither rocher:et:al:2019 nor dick:etal:2023 propose the leave-one-out approach to estimating the extent of confidentiality violation we pursue here, though dick:etal:2023 analyze an interesting alternative set of baselines.
2.5. Statistical inference vs. breach of confidentiality
@@
Privacy and confidentiality protections are subtle concepts that give rise to many misunderstandings. It is a common but incorrect belief that breaches of confidentiality, colloquially known as privacy breaches, occur when a dataset is used to make any harmful or unwanted inference about an individual. Such an error has made its way into many peer-reviewed papers (jarmin:etal:2023, SI section 5). In order to see where the problem lies, we first discuss the canonical “smoking causes cancer” thought experiment (dwork:2011) and then discuss possible confidentiality concerns in the 2020 Census data. A more complete version of these arguments can be found in kifer:etal:2022.
@@
The first Cancer Prevention Study, also known as CPS-I, followed a cohort of volunteers from 1959 to 1972 and conclusively established the link between smoking tobacco cigarettes as a cause of death from lung cancer and coronary heart disease (cps1). As a result of this study, we know that persons who smoke have a much higher risk of developing lung cancer. Such inferences about smokers can definitely be unwanted, as they result in higher health and life insurance premiums. Persons born after 1972 may be subject to this inference caused by the study; however, since their data were not used for the study (they were not born until after it was completed), the study cannot possibly be considered a confidentiality breach of their data. For those people, one would say that the inference is purely statistical in nature. Another way to phrase this is that the link between smoking and lung cancer is a population property, or statistical use (in the sense of the 2018 Confidential Information Protection and Statistical Efficiency Act, Title III of the 2018 Foundations of Evidence-based Policymaking Act; 44 U.S. Code § 3561(12)) that was uncovered with the help of the data set. Enabling generalizable scientific inference like this, or inference about large-scale populations patterns, is exactly why data sets are collected and published.
@@
In contrast, an unwanted inference is a confidentiality breach when it is specifically caused by the inclusion of the individual’s information in the dataset from which the inference was made, what we call non-LOO inference. Now consider a hypothetical CPS-I study participant named Charlie who was a lifelong smoker. Suppose that as a result of the study, Charlie’s insurance company decided to ask enrollees whether or not they smoked and charge a higher premium if they did. As a result, Charlie was harmed by the result of the study in which he was a participant. Is this now a confidentiality breach? To answer that question, we turn to causal reasoning, and specifically consider a counterfactual in which Charlie had not participated in the study; that is, we compare the non-LOO inference with a properly computed LOO one. Would the outcome of the study have been different enough without Charlie’s participation to change the findings—thus changing whether the insurance company enacted their policy of charging smokers a higher premium policy which harmed Charlie? Given the strength of the findings in CPS-I (hammond:horn:1954), we can be confident that the study would have drawn the same conclusions even if Charlie had not participated. Therefore, this example would also not be considered a confidentiality breach.
@@
One also must be careful about the conflation of harmful inferences with confidentiality violations for another reason, which is also related to privacy and trust in statistical agencies, again as they are expressed in 44 U.S. Code § 3561-4 and nasem:2021. Specifically, sometimes a statistical inference should not be allowed even if it is not confidentiality violating. A statistical agency might be asked to produce data on a particular sensitive population that could reasonably expect harm from those data even if they passed the LOO inference test. For example, in 2010, communities with a high percentage of same-sex married couples in states that did not permit such weddings might expect harm even if only a statistical summary were published. A statistical summary that could have affected the 2020 Census is the proportion of residents in the census block who were not U.S. citizens. Whether or not the government has a legitimate statistical interest in these populations is indeed a policy question, but the policy concerns ingesting the data with the intention to publish summaries in the first place. Barring the collection or publication of data on the grounds that even statistical inference may be harmful is a policy concern. In this paper, we presume that the agency’s legitimate interest in supporting statistical inference has already been determined; that is, the collection of data in the decennial census and their publication for statistical purposes are authorized by Congress and undertaken consistent with the required trust and confidentiality policies.
In practice, the distinction between confidentiality-violating and statistical inferences is not always so clear-cut as in our smoking example. Information obtained from individuals is often aggregated, fields are suppressed (e.g., direct identifiers), and some SDL framework beyond aggregation may also be used. Still, it is important to disentangle what could be learned from statistical inference versus what can be specifically learned or caused by a person’s participation/inclusion in a study or dataset. This idea is the basis for research in the scientific field of “differential privacy” (Dwork:Algorithmic:Book:2014).
3. Overview of Workflow, 2010 Census, and Commercial Data Inputs